m^4-n^4

8b²+24b+18
3x²­18x-48
9x²+4y²




Answer by stanbon(75887)

m^4-n^4
= (m^2+n^2)(m^2-n^2)
= (m^2+n^2)(m+n)(m-n)
---------------------------
8b²+24b+18
= 2(4b^2 + 12b + 9)
= 2(4b^2 + 6b + 6b + 9)
= 2(2b(2b+3) + 3(2b+3))
= 2(2b+3)(2b+3)
=====================
3x²+ ­18x - 48
= (x-2)(3x+24)
=====================
9x²+4y²
not factorable in the Real Number system.
-----
(3x+2iy)(3x-2iy) in the Complex number system.
======================
Cheers,
Stan H.